Two different types of drugs A and B were tried on different patients for increasing weight. 5 persons were given A and 7 persons were given B. Increase in weights(in. kg) are given below:
A: 3.6 5.5 5.9 4.1 1.4
B: 4.5 3.6 5.5 6.8 2.7 3.6 5.0
Do the drugs differ significantly with regard to their effect in increasing weight?

----- This question is from the Test of Hypothesis -----
1
Expert's answer
2020-09-20T17:55:34-0400

Using TI84, the provided sample means are shown below:

"\bar X_1 = 4.1" "\bar X_2" =4.529

Also, the provided sample standard deviations are :"s \n_1\n\u200b\t\n =1.785;\ns_2 = 1.378"

and the sample sizes are  "n_1 = 5\n\nand" "n_2 = 7"


The following null and alternative hypotheses need to be tested:

"Ho: \mu_1= \mu_2"

"Ha: \mu_1≠ \mu_2"


This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

Based on the information provided, the significance level is α=0.05, and the degrees of freedom are df = "df=\frac{\frac{s_{1}^{2}}{n_1}+\frac{s_{2}^{2}}{n_2}}{\frac{\frac{s_{1}^{2}}{n_1}}{n_1-1}+\frac{\frac{s_{2}^{2}}{n_2}}{n_2-1}}" ="df=\frac{\frac{1.785^{2}}{5}+\frac{1.378^{2}}7}{\frac{\frac{1.785^{2}}5}{5-1}+\frac{\frac{1.378^{2}}{7}}{7-1}}"= 7.254

Since it is assumed that the population variances are unequal, the t-statistic is computed as follows:

"t = \frac{\overline{X1}-\overline{X2}}{\sqrt{\frac{(s1^2)}{n1}+\frac{s2^2}{n2}}}" = "\frac{4.1-4..529}{\sqrt{\frac{(1.7851^2)}{5}+\frac{1.378^2}{7}}}" = -0.45

Hence, it is found that the critical value for this two-tailed test is tc

​=2.348, for α=0.05 and df = 7.254, using t table the rejection region for this two-tailed test is R={t:∣t∣>2.348}.since critical value is greater than test statistic ,it is concluded that the null hypothesis is not rejected.

t is concluded that the null hypothesis Ho is not rejected. Therefore, there is not enough evidence to claim that the population mean μ1 is different than μ2​, at the 0.05 significance level.